how to solve -5(4x+8)=-260
Answer by oberobic(2304)

The short answer is "methodically."
.
-5(4x+8)=-260
.
divide both sides by -5 to get rid of it on the left-hand side, and it will make the right-hand side positive.
(4x+8) = 52
4x + 8 = 52
.
subtract 8 from both sides
4x = 44
.
divide both sides by 4
x = 11
.
check by substituting back into the original equation
.
-5(4x+8)=-260 ??
-5(4*11+8) = -5(52) = -260
correct.
.
Answer:
x = 11
.
Done.